OpenStudy (debbielve):

Can someone please work this out for me I am not understanding it. If 𝑓(𝑥) = cos 𝜋𝑥, find the exact value of a such that 𝑓(1)−𝑓(−1)/1-(-1)= cos a.

jimthompson5910 (jim_thompson5910):

Plug in x = 1 to get \[\Large f(x) = \cos(\pi*x)\] \[\Large f(1) = \cos(\pi*1)\] \[\Large f(1) = \cos(\pi)\] \[\Large f(1) = -1\] You'll use the unit circle https://upload.wikimedia.org/wikipedia/commons/thumb/4/4c/Unit_circle_angles_color.svg/2000px-Unit_circle_angles_color.svg.png to compute the last part. What do you get when you plug in x = -1 ?

OpenStudy (sshayer):

\[\frac{ \cos \pi-\cos (-\pi) }{ 1-(-1) }=\frac{ \cos \pi-\cos \pi }{ 2 }=0=\cos a\] \[\cos a=0=\cos \left( 2n +1 \right)\frac{ \pi }{ 2 },\] where n is an integer.

OpenStudy (sshayer):

\[\cos (-x)=\cos x\]

OpenStudy (sshayer):

a=?